OpenStudy (anonymous):

how would i simplify x2+2x+1 ------- -1-x

OpenStudy (anonymous):

using -1 as a factor

OpenStudy (anonymous):

Assuming your expression is\[\frac{x^2+2x+1}{-1-x}\]You can re-write it as\[\frac{(x+1)^2}{-(x+1)}=-\frac{(x+1)^2}{(x+1)}\]The quadratic in the numerator may be factored as (x+1)^2 and numerator as -(x+1). Notice there is a common factor of (x+1) in the numerator and denominator. These cancel to 1, and you're left with,\[-(x+1)\]Hope this helps :)